OpenStudy (anonymous):

Add and simplify: x+1/x^2+3/4x 11/4x, x ≠ 0 4x+7/4x^2, x ≠ 0 x+4/4x^2, x ≠ 0 7x+4/4x^2, x ≠ 0

OpenStudy (anonymous):

help didn't understand the lesson in need of assistants!!

OpenStudy (anonymous):

I would help but I don't even get the question. What was the lesson about?

OpenStudy (anonymous):

it is about adding and subtracting rational expressions

OpenStudy (anonymous):

thanks for trying though

OpenStudy (anonymous):

Do you mean \[\frac{ x+1 }{ x^2 } + \frac{ 3 }{ 4x }\]

OpenStudy (anonymous):

for teh first one?

OpenStudy (anonymous):

ya for the question

OpenStudy (anonymous):

oh then maybe i can help after all

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

my online teacher wasn't helping me so i really am confused on this lesson

OpenStudy (anonymous):

if you multiply \[\frac{ 3 }{ 4x }\] by x, you get \[\frac{ 3x }{ 4x^2 }\] which will simplify your question to \[\frac{ x+1 }{ x^2 } + \frac{ 3x }{ 4x^2 }\] (since \[\frac{ 3 }{ 4x } = \frac{ 3x }{ 4x^2 }\] the x cancel out on the right hand side) we have to now multiply the top and bottom of \[\frac{ x+1}{ x^2 }\] by 4 to get the same denominator. This will give us \[\frac{ 4(x+1) }{ 4x^2 }\] Now the question becomes \[\frac{ 4(x+1) }{ 4x^2 }\ + \frac{ 3x }{ 4x^2 }\] which is equals to \[\frac{ 4(x+1) +3x }{ 4x^2 }\] which gives us the answer \[\frac{ 7x+4 }{ 4x^2 }\] Hope you get it...